OpenStudy (anonymous):

please help me on this one: A leather coin purse has the shape of a triangle. Two sides are equal in length and the third side is 3 centimeters shorter than on and one half times the length of the equal sides. The perimeter is 28.5 centimeters. Find the lengths of the sides.

OpenStudy (anonymous):

\[perimeter=\sum_{}^{} legs\] \[28.5=x+x +x/2\] \[x=11.4\] \[l _{1}=l_{2}=11.5cm\] \[l_{3}=11.5/2=5.75\]

OpenStudy (anonymous):

\[perimeter=\sum_{}^{} legs\] \[28.5=x+x +x/2\] \[x=11.4\] \[l _{1}=l_{2}=11.5cm\] \[l_{3}=11.5/2=5.75\]